the length of a rectangle is 2 more than 3 times the width. if the perimeter is 36 feet, find the dimensions of the rectangle
Answer by TimothyLamb(4379)

L = 3w + 2
2L + 2w = 36
---
2L + 2w = 36
2(3w + 2) + 2w = 36
6w + 4 + 2w = 36
8w = 32
---
w = 4
L = 3w + 2
L = 12 + 2
L = 14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php